### v1

@@ Abstract
Auto-SPICE is the first fully automated framework leveraging large language models (LLMs) to generate Simulation Programs with Integrated Circuit Emphasis (SPICE) netlists. It addresses a long-standing challenge in automating netlist generation for analog circuits within circuit design automation. Automating this workflow could accelerate the creation of finetuned LLMs for analog circuit design and verification. We identify key challenges in this automation and evaluate the multi-modal capabilities of state-of-the-art LLMs, particularly GPT-4, to address these issues. We propose a three-step workflow to overcome current limitations: labeling analog circuits, prompt tuning, and netlist verification. This approach aims to create an end-to-end SPICE netlist generator from circuit schematic images, tackling the long-standing hurdle of accurate netlist generation. Our framework demonstrates significant performance improvements, tested on approximately 2,100 schematics of varying complexity. We open-source this solution for community-driven development.
@@ I Introduction
Contributions: Auto-SPICE is an automated method for SPICE netlist generation from analog circuit schematics. We incorporate various techniques, including (i) fine-tuned object detectors to extract passive and active components, (ii) deep Hough transform priors to identify nets, (iii) extensive prompt tuning for LLMs to fix common errors in netlist extraction, and (iv) post-extraction verification. We make four contributions:
@@
    Build Auto-SPICE, the first fully automated framework using LLMs along with custom-trained deep network models for large-scale SPICE netlist extraction from circuit schematics present in textbooks and papers.
@@
    Use Auto-SPICE to collect a dataset of $\sim$2,100 SPICE netlists (along with metadata like figure captions) from a sample textbook. We open-source our flow to the community to further research in this area [16].
@@
    Fine-tune GPT models using the Auto-SPICE dataset to demonstrate the potential use-case of the dataset in automated SPICE netlist generation from English prompts.
@@ IV Methodology
Figure 4: Dataset preparation flow starting from a PDF document consisting of Schematic and related text. Auto-SPICE extracts the relevant information and generates the SPICE netlist for the schematic.
@@
This section details the methodology developed in this study, addressing the significant challenges outlined in Section III. Specifically, the accurate generation of SPICE netlists remains a critical bottleneck for LLMs due to difficulties in correctly identifying circuit components and establishing net connections. To begin the dataset creation flow, as shown in Fig. 4 (➊), it takes a document in PDF format. After that, it is subdivided into 2 parts: (I) Extraction/Summarization of text present in the document comprises of the description of schematic and Key Performance Indicators (KPI) like gain, linearity, bandwidth, noise, gain/phase margin, etc - This will be used as an additional context for fine-tuning LLMs, Section V-D, shown in Fig. 4 (➏ - ➐); (II) Automatic SPICE Netlist Generation - Auto-SPICE: This step starts with extraction of the schematics/image from the document, then goes through the three-step process: 1) Labeling Analog Circuit, 2) Prompt Tuning, and 3) SPICE netlist verification, to finally generate a SPICE netlist, shown in Fig. 4 (➋ - ➎).
@@ V-A Dataset Creation
Figure 7: Automatic SPICE Netlist generation using Auto-SPICE.
@@
We curated a comprehensive dataset by selecting schematics from textbooks [24, 30]. Textbooks were chosen over research papers because they offer high-quality schematic images, clear context, and detailed explanations. This makes them more suitable for our purposes, as research papers often present high-level descriptions with advanced schematics that use abstract blocks for complex components. We collected over 2,100 schematic images and applied our proposed Auto-SPICE flow (Fig. 4) to generate the corresponding SPICE netlists. Alongside this, we collected captions and descriptions as detailed in Section IV. Fig. 7 demonstrates how Auto-SPICE translates a schematic into a SPICE netlist. Since our primary goal is to generate structurally correct netlists, each component is marked with default parameters.
@@ V-C Verifying against AMSNet [13] netlists
We demonstrated the effectiveness of our approach by evaluating it using the AMSNet [13] dataset, which includes $\sim$800 circuit diagrams. Unlike AMSNet, which relies on an algorithmic method supplemented with some manual effort to generate SPICE netlists, our method introduces a fully automated flow (Fig. 4). When comparing the SPICE netlists generated by our method to those produced by AMSNet, our approach achieved a 100% similarity score, as defined earlier. This validates the robustness of our flow and motivates us to extend the evaluation to additional circuit diagrams. Fig. 9 illustrates an example of graphs representation of SPICE netlist generated by AMSNet (➊) and our Auto-SPICE framework (➋). The results demonstrate that the number of components and their neighboring elements is identical across both methods, with differences in component and net naming conventions. These findings strengthen Auto-SPICE accuracy and scalability in generating accurate netlists for various circuit diagrams.
Figure 9: Graph representation of SPICE netlists generated by AMSNet [13] and Auto-SPICE. They are identical.
@@ V-D Finetuning
Table I reports preliminary results of fine-tuning the GPT’s models on the Auto-SPICE datasets generated. We picked 20 common design problems frequently used as a standalone circuit or sub-part of a bigger design. We prompted the LLM to generate (n = 10) samples of SPICE netlist for each circuit description and manually validated their correctness (the score reported is the number of correct SPICE netlists). We subdivided the 20 benchmarks into 3 different categories: Easy (ID=1-7), Medium (ID=8-14), and Hard (ID=15-20).  Table I provides a comprehensive evaluation of the performance of GPT-3.5-turbo and GPT-4o-mini, both in their ‘base’ and ‘finetune’ states, on a diverse set of circuit design tasks. Finetuning demonstrates significant performance gains across nearly all circuits, particularly for complex designs. For instance, in the (ID=8), GPT-3.5-turbo improves from a score of 4 to 10. At the same time, GPT-4o-mini achieves a perfect score of 10 after fine-tuning, demonstrating the impact of domain-specific training. This pattern is evident across other tasks such as the (ID=5), where fine-tuning boosts GPT-3.5-turbo from 5 to 9 and GPT-4o-mini from 6 to 10.
@@ VI Conclusion and Future Work
In this paper, we introduced Auto-SPICE, a novel framework to create large-scale datasets for SPICE netlists from documents comprising analog circuit schematics, addressing critical challenges in the design workflow. Through schematic preprocessing and prompt tuning, we developed a fully automated pipeline that leverages LLMs for scalable netlist generation. To the best of our knowledge, Auto-SPICE is the first LLM-based framework of its kind, and our released dataset and flow [16] provide a valuable resource for the community for future advancements.

### v2

@@ Abstract
Masala-CHAI is the first fully automated framework leveraging large language models (LLMs) to generate Simulation Programs with Integrated Circuit Emphasis (SPICE) netlists. It addresses a long-standing challenge in automating netlist generation for analog circuits within circuit design automation. Automating this workflow could accelerate the creation of finetuned LLMs for analog circuit design and verification. We identify key challenges in this automation and evaluate the multi-modal capabilities of state-of-the-art LLMs, particularly GPT-4, to address these issues. We propose a three-step workflow to overcome current limitations: labeling analog circuits, prompt tuning, and netlist verification. This approach aims to create an end-to-end SPICE netlist generator from circuit schematic images, tackling the long-standing hurdle of accurate netlist generation. Our framework demonstrates significant performance improvements, tested on approximately 2,100 schematics of varying complexity. We open-source this solution for community-driven development.
@@ I Introduction
Contributions: Masala-CHAI [^1] Masala is the Hindi word for “SPICE” is an automated method for SPICE netlist generation from analog circuit schematics. We incorporate various techniques, including (i) fine-tuned object detectors to extract passive and active components, (ii) deep Hough transform priors to identify nets, (iii) extensive prompt tuning for LLMs to fix common errors in netlist extraction, and (iv) post-extraction verification. We make four contributions:
@@
    Build Masala-CHAI, the first fully automated framework using LLMs along with custom-trained deep network models for large-scale SPICE netlist extraction from circuit schematics present in textbooks and papers.
@@
    Use Masala-CHAI to collect a dataset of $\sim$2,100 SPICE netlists (along with metadata like figure captions) from a sample textbook. We open-source our flow to the community to further research in this area [16].
@@
    Fine-tune GPT models using the Masala-CHAI dataset to demonstrate the potential use-case of the dataset in automated SPICE netlist generation from English prompts.
@@ IV Methodology
Figure 4: Dataset preparation flow starting from a PDF document consisting of Schematic and related text. Masala-CHAI extracts the relevant information and generates the SPICE netlist for the schematic.
@@
This section details the methodology developed in this study, addressing the significant challenges outlined in Section III. Specifically, the accurate generation of SPICE netlists remains a critical bottleneck for LLMs due to difficulties in correctly identifying circuit components and establishing net connections. To begin the dataset creation flow, as shown in Fig. 4 (➊), it takes a document in PDF format. After that, it is subdivided into 2 parts: (I) Extraction/Summarization of text present in the document comprises of the description of schematic and Key Performance Indicators (KPI) like gain, linearity, bandwidth, noise, gain/phase margin, etc - This will be used as an additional context for fine-tuning LLMs, Section V-D, shown in Fig. 4 (➏ - ➐); (II) Automatic SPICE Netlist Generation - Masala-CHAI: This step starts with extraction of the schematics/image from the document, then goes through the three-step process: 1) Labeling Analog Circuit, 2) Prompt Tuning, and 3) SPICE netlist verification, to finally generate a SPICE netlist, shown in Fig. 4 (➋ - ➎).
@@ V-A Dataset Creation
Figure 7: Automatic SPICE Netlist generation using Masala-CHAI.
@@
We curated a comprehensive dataset by selecting schematics from textbooks [24, 30]. Textbooks were chosen over research papers because they offer high-quality schematic images, clear context, and detailed explanations. This makes them more suitable for our purposes, as research papers often present high-level descriptions with advanced schematics that use abstract blocks for complex components. We collected over 2,100 schematic images and applied our proposed Masala-CHAI flow (Fig. 4) to generate the corresponding SPICE netlists. Alongside this, we collected captions and descriptions as detailed in Section IV. Fig. 7 demonstrates how Masala-CHAI translates a schematic into a SPICE netlist. Since our primary goal is to generate structurally correct netlists, each component is marked with default parameters.
@@ V-C Verifying against AMSNet [13] netlists
We demonstrated the effectiveness of our approach by evaluating it using the AMSNet [13] dataset, which includes $\sim$800 circuit diagrams. Unlike AMSNet, which relies on an algorithmic method supplemented with some manual effort to generate SPICE netlists, our method introduces a fully automated flow (Fig. 4). When comparing the SPICE netlists generated by our method to those produced by AMSNet, our approach achieved a 100% similarity score, as defined earlier. This validates the robustness of our flow and motivates us to extend the evaluation to additional circuit diagrams. Fig. 9 illustrates an example of graphs representation of SPICE netlist generated by AMSNet (➊) and our Masala-CHAI framework (➋). The results demonstrate that the number of components and their neighboring elements is identical across both methods, with differences in component and net naming conventions. These findings strengthen Masala-CHAI accuracy and scalability in generating accurate netlists for various circuit diagrams.
Figure 9: Graph representation of SPICE netlists generated by AMSNet [13] and Masala-CHAI. They are identical.
@@ V-D Finetuning
Table I reports preliminary results of fine-tuning the GPT’s models on the Masala-CHAI datasets generated. We picked 20 common design problems frequently used as a standalone circuit or sub-part of a bigger design. We prompted the LLM to generate (n = 10) samples of SPICE netlist for each circuit description and manually validated their correctness (the score reported is the number of correct SPICE netlists). We subdivided the 20 benchmarks into 3 different categories: Easy (ID=1-7), Medium (ID=8-14), and Hard (ID=15-20).  Table I provides a comprehensive evaluation of the performance of GPT-3.5-turbo and GPT-4o-mini, both in their ‘base’ and ‘finetune’ states, on a diverse set of circuit design tasks. Finetuning demonstrates significant performance gains across nearly all circuits, particularly for complex designs. For instance, in the (ID=8), GPT-3.5-turbo improves from a score of 4 to 10. At the same time, GPT-4o-mini achieves a perfect score of 10 after fine-tuning, demonstrating the impact of domain-specific training. This pattern is evident across other tasks such as the (ID=5), where fine-tuning boosts GPT-3.5-turbo from 5 to 9 and GPT-4o-mini from 6 to 10.
@@ VI Conclusion and Future Work
In this paper, we introduced Masala-CHAI, a novel framework to create large-scale datasets for SPICE netlists from documents comprising analog circuit schematics, addressing critical challenges in the design workflow. Through schematic preprocessing and prompt tuning, we developed a fully automated pipeline that leverages LLMs for scalable netlist generation. To the best of our knowledge, Masala-CHAI is the first LLM-based framework of its kind, and our released dataset and flow [16] provide a valuable resource for the community for future advancements.